OpenStudy (anonymous):

a barrel contains 187 gallons of water and is being filled at a contant rate of 8 gallons per hour. write an equation that models the number of gallons.g after t hours. the equation that models the given situation is____.

OpenStudy (zzr0ck3r):

we start with 187 gallons, and we are filling at 8g/h example: so if we fill for 8 seconds we add 8(g/h)*8*h = 64g then we add that to 187 for our total so f(t) = 187 + 8t

OpenStudy (anonymous):

ok im lost

OpenStudy (zzr0ck3r):

imagine you had 20 bucks and I gave you another 5 bucks an hour then the function would be f(x) = 20+5t does this makes sense?

OpenStudy (zzr0ck3r):

just like that you answer is f(x) = 187 + 8t